OpenStudy (hunter1222):

An electric device uses 650 watts of power. If the voltage is 120 volts, what is the resistance?

OpenStudy (salam):

22.25 ohm

OpenStudy (radar):

Use the equation P =(E^2)/R. Substitute the given values 650 = (120)(120)/R R = (120X120)/650 R = (24X12)/13 R = 288/13 Ohms R = 22.1538 Ohms

OpenStudy (skybrel):

Hello, simply first we calculate Current in Amperes. I (current)= P (power in watts) / Voltage(in volts ) I= 650/120 , in Ampere I=5.416 A (Ampere) now Applying Ohm's Law V=IR we have V=120V and I=5.416A and R=? R=V/I , So, R=120/5.416=22.156573 so R=22.156573 ohms [where ohms represent by symbol Ω].. So Finally R=22.156573Ω Thanks.

OpenStudy (radar):

Note that their are 3 results posted. The difference is in the manner in which "rounding off" was accomplished. The value of 22.1538 Ohms was obtained by keeping the answer in fractional form until the final calculation.